A study was conducted to investigate the relationship between roadway surface temperature (x) and pavement deflection (y). The data are as follows:


x

65.3

66.4

70.8

70.7

69.8

71.6

72.1

72.4

73.5


y

0.6144

0.6273

0.6668

0.6781

0.6497

0.6678

0.6783

0.6755

0.6873


Calculate the Pearson correlation coefficient. Round your answers to the nearest hundredths.


1
Expert's answer
2022-05-31T10:58:20-0400

The following table needs to be used:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 X & Y & XY & X^2 & Y^2 \\ \hline
65.3 & 0.6144 & 40.12032 & 44264.09 & 0.37748736 \\
 \hdashline
66.4 & 0.6273 & 41.65272 & 4408.96 & 0.39350529 \\
 \hdashline
 70.8 & 0.6668 & 47.20944 & 5012.64 & 0.44462224 \\
 \hdashline
70.7 & 0.6781 & 47.94167 & 4998.49 & 0.45981961 \\
 \hdashline
69.8 & 0.6497 & 45.34906 & 4872.04 & 0.42211009 \\
 \hdashline
71.6 & 0.6678 & 47.81448 & 5126.56 & 0.44595684 \\
 \hdashline
72.1 & 0.6783 & 48.90543 & 5198.41 & 0.46009089 \\
 \hdashline
72.4 & 0.6755 & 48.9062 & 5241.76 & 0.45630025 \\
 \hdashline
73.5 & 0.6873 & 50.51655 & 5402.25 & 0.47238129 \\
 \hdashline
Sum =\\
632.6 & 5.9452 & 418.41587 & 44525.2 & 3.93227386\\
 \hdashline
\end{array}"




"\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{632.6}{9}""=70.288889""\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{5.9452}{9}""=0.660578""SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=44525.2-\dfrac{632.6^2}{9}=60.448889""SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=3.93227386-\dfrac{5.9452^2}{9}=0.005007"




"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=418.41587-\dfrac{632.6(5.9452)}{9}=0.534368""r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}SS_{YY}}}=\dfrac{0.534368}{\sqrt{60.448889(0.005007)}}"




"=0.97132277"